OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $8 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $194 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 194 = one-halfx + 2(x - 8) What is the regular price of a child's ticket? $68 $84 $89 $105

OpenStudy (anonymous):

You'll want to start by trying to isolate x. Start by distributing the 2 onto (x-8) 194 = 0.5x + 2x - 16 Now you can combine like terms 194 = 2.5x - 16 Add 16 to both sides 194 + 16 = 2.5x Now divide both sides by 2.5 to get x by itself (194 + 16) / 2.5 = x 210 / 2.5 = x x will be the price of an adult ticket. The answer is $84. Now to find the price of a child's ticket, just subtract 8 from that, which is $76. So your answer is B

OpenStudy (anonymous):

thanks(: